Constants Periodic Table A 1.0-L buffer solution initially contains 0.30 mol of NH: (K) = 1.76 x 10-5) and 0.30 mol of NH CI.Part A)

In order to adjust the buffer pH to 8.65, should you add NaOH or HCl to the buffer mixture?

- Correct Answer: HCL

Part B)

What mass of the correct reagent should you add?

Express your answer using two significant figures.
Answer #1

The pOH of basic buffer is given by pOH = pKb + log(NH4Cl/NH3)

                                                                = -log(1.76*10^-5) + log(0.30/0.30)

The pOH of given basic buffer is 4.75

In order to adjust the buffer pH to 8.65, we should add HCl to the buffer mixture

If x is the moles of HCl added to the buffer solution, it reacts with NH3 to form NH4Cl so that there is a decrease in the concentration of the NH3 by x moles and increase in the concentration of NH4Cl by x moles.

pH is given as 8.65. ======> pOH = 14-8.65 =5.35

pOH = pKb + log(NH4Cl/NH3)

5.35 = -log(1.76*10^-5) + log(0.30+x/0.30-x)

number of moles of HCl added (x) = 0.178 moles

The mass of the correct reagent(HCl) should be added =0.178*36.5 = 6.5 grams